OpenStudy (anonymous):

Solve x + 7 = –2(4x + 10). You must show all work.

OpenStudy (anonymous):

x+7=-8x-20 Now x+8x=-20-7 Implies 9x=-27 Now x= -27/9 = -3 That is x= -3

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

Questions?